# Verifiable type-III seesaw and dark matter in a gauged $\boldsymbol{U(1)_{\rm B-L}}$ symmetric model

### v2

@@ Appendix A Triplet fermion decay modes
The electroweak correction induces a small mass splitting between the charged and neutral components of the triplet. This mass splitting is given by [32]:
@@
The mixing between the leptons and heavy triplet fermion states can be expressed as [27]
@@
By choosing $\mathcal{R}$ to be an identity matrix it has been shown that the above decay modes (Eq. 49-54) of the 1st generation of triplet are sensitive to the lightest neutrino mass ($m_{1}$ for NH and $m_{3}$ for IH) while the decay modes of 2nd and 3rd generation of triplets are independent of lightest neutrino mass [27].
@@ Appendix C Gravitational waves from FOPT
$f^{{\rm col}}_{peak}$ $=$ $1.65\times 10^{-5}({\rm Hz})\bigg(\frac{g_{*}(T_{n})}{100}\bigg)^{1/6}\bigg(\frac{T_{n}}{0.1{~\rm TeV}}\bigg)$ (70)
@@
$\bigg(\frac{\kappa_{\phi}(T_{n})\alpha(T_{n})}{1+\alpha(T_{n})}\bigg)^{2}\frac{A(a+b)^{c}}{\bigg(b\big(\frac{f}{f^{{\rm col}}_{peak}}\big)^{-a/c}+a\big(\frac{f}{f^{{\rm col}}_{peak}}\big)^{b/c}\bigg)^{c}},$
@@
The peak frequency and the amplitude of the GWs generated by the sound waves are expressed by [96]
$f^{{\rm sw}}_{peak}$ $=$ $8.9\times 10^{-6}({\rm Hz})\bigg(\frac{g_{*}(T_{n})}{100}\bigg)^{1/6}\bigg(\frac{1}{v_{w}}\bigg)\bigg(\frac{T_{n}}{0.1{~\rm TeV}}\bigg)$ (73)
@@
$\bigg(\frac{\kappa_{\rm sw}(T_{n})\alpha(T_{n})}{1+\alpha(T_{n})}\bigg)^{2}\frac{7^{3.5}\big(\frac{f}{f^{{\rm sw}}_{peak}(T_{n})}\big)^{3}}{\bigg(4+3\big(\frac{f}{f^{{\rm col}}_{peak}(T_{n})}\big)^{2}\bigg)^{3.5}}.$
The efficiency factor is [98]
@@
where the lifetime of the sound wave is given as [99]
@@
The peak frequency and the amplitude of the GWs generated by the turbulence are expressed by [100, 96, 97]
$f^{{\rm turb}}_{peak}$ $=$ $2.7\times 10^{-5}({\rm Hz})\bigg(\frac{g_{*}(T_{n})}{100}\bigg)^{1/6}\bigg(\frac{1}{v_{w}}\bigg)\bigg(\frac{T_{n}}{0.1{~\rm TeV}}\bigg)$ (80)
@@
$v_{w}\frac{\big(\frac{f}{f^{{\rm sw}}_{peak}(T_{n})}\big)^{3}}{\bigg(1+\big(\frac{f}{f^{{\rm col}}_{peak}(T_{n})}\big)\bigg)^{3.6}\big(1+\frac{8\pi f}{h_{*}(T_{n})}\big)},$
where the efficiency factor is $\kappa_{{\rm turb}}=0.1\kappa_{{\rm sw}}$ [100], and the inverse of the Hubble time at the epoch of GW production, red-shifted to today, is expressed as:

### v3

@@ Appendix A Triplet fermion decay modes
The electroweak correction induces a small mass splitting between the charged and neutral components of the triplet. This mass splitting is given by Cirelli:2005uq :
@@
The mixing between the leptons and heavy triplet fermion states can be expressed as Das:2020uer
@@
By choosing $\mathcal{R}$ to be an identity matrix it has been shown that the above decay modes (Eq. 49-54) of the 1st generation of triplet are sensitive to the lightest neutrino mass ($m_{1}$ for NH and $m_{3}$ for IH) while the decay modes of 2nd and 3rd generation of triplets are independent of lightest neutrino mass Das:2020uer .
@@ Appendix C Gravitational waves from FOPT
$f^{{\rm col}}_{\rm peak}$ $=$ $1.65\times 10^{-5}({\rm Hz})\bigg(\frac{g_{*}(T_{n})}{100}\bigg)^{1/6}\bigg(\frac{T_{n}}{0.1{~\rm TeV}}\bigg)$ (70)
@@
$\bigg(\frac{\kappa_{\phi}(T_{n})\alpha(T_{n})}{1+\alpha(T_{n})}\bigg)^{2}\frac{A(a+b)^{c}}{\bigg(b\big(\frac{f}{f^{{\rm col}}_{\rm peak}}\big)^{-a/c}+a\big(\frac{f}{f^{{\rm col}}_{\rm peak}}\big)^{b/c}\bigg)^{c}},$
@@
The peak frequency and the amplitude of the GWs generated by the sound waves are expressed by Athron:2023xlk
$f^{{\rm sw}}_{\rm peak}$ $=$ $8.9\times 10^{-6}({\rm Hz})\bigg(\frac{g_{*}(T_{n})}{100}\bigg)^{1/6}\bigg(\frac{1}{v_{w}}\bigg)\bigg(\frac{T_{n}}{0.1{~\rm TeV}}\bigg)$ (73)
@@
$\bigg(\frac{\kappa_{\rm sw}(T_{n})\alpha(T_{n})}{1+\alpha(T_{n})}\bigg)^{2}\frac{7^{3.5}\big(\frac{f}{f^{{\rm sw}}_{\rm peak}(T_{n})}\big)^{3}}{\bigg(4+3\big(\frac{f}{f^{{\rm col}}_{\rm peak}(T_{n})}\big)^{2}\bigg)^{3.5}}.$
The efficiency factor is Espinosa:2010hh
@@
where the lifetime of the sound wave is given as Guo:2020grp
@@
The peak frequency and the amplitude of the GWs generated by the turbulence are expressed by Caprini:2015zlo ; Athron:2023xlk ; Caprini:2024hue
$f^{{\rm turb}}_{\rm peak}$ $=$ $2.7\times 10^{-5}({\rm Hz})\bigg(\frac{g_{*}(T_{n})}{100}\bigg)^{1/6}\bigg(\frac{1}{v_{w}}\bigg)\bigg(\frac{T_{n}}{0.1{~\rm TeV}}\bigg)$ (80)
@@
$v_{w}\frac{\big(\frac{f}{f^{{\rm sw}}_{\rm peak}(T_{n})}\big)^{3}}{\bigg(1+\big(\frac{f}{f^{{\rm col}}_{\rm peak}(T_{n})}\big)\bigg)^{3.6}\big(1+\frac{8\pi f}{h_{*}(T_{n})}\big)},$
where the efficiency factor is $\kappa_{{\rm turb}}=0.1\kappa_{{\rm sw}}$ Caprini:2015zlo , and the inverse of the Hubble time at the epoch of GW production, red-shifted to today, is expressed as: